OpenStudy (anonymous):

Write an equation of the line with the given slope and y-intercept slope: ,y-intercept: –9 a.y = x– 9 b.y = x – 9 c.y = x – 9 d.y = x+ 9

OpenStudy (anonymous):

@KamiBug

OpenStudy (solomonzelman):

I can see from the choices that the slope is 1. So, let's rephrase your question \(\color{red}{Write~~~~an~~~~equation~~~~of~~~~the ~~~~line~~~~ with ~~~~the~~~~ given~~~~ slope ~~~~and }\) \(\color{red}{y-intercept. }\) \(\color{blue}{slope :~ 1~~~~~~~~~~~~~~y-intercept:~~9 }\) the form is \(\color{blue}{y=mx+b }\) where m is the slope and b is the y-intercept plug the numbers in \(\color{red}{y=\color{blue}{1}~~ x+\color{blue}{-9}}\) simplify that, and you get \[\color{blue}{y=x-9}\]

OpenStudy (solomonzelman):

I made a typo y-intercept is -9, sorry :)

OpenStudy (anonymous):

@SolomonZelman slope: 1.6, y-intercept: 7.5 a. y = 1.6x – 7.5 c. y = 1.6x + 7.5 b. y = 646-06-03-00-00_files/i0150000.jpgx + 7.5 d. y = –1.6x + 7.5

OpenStudy (solomonzelman):

plug them in y= (1.6)x+(7.5) and that's the same as, y=1.6x+7.5

OpenStudy (kamibug):

So sorry I'm late! I was trying to come but my computer hates me. Anyways, @SolomonZelman is correct in both cases! It's just all about substitution! =) I promise next time I'll try to be here asap, LOL!!! xD Okk, sorry.